OpenStudy (anonymous):

please, one more? I'll help you in any other subject.... In the below system, solve for y in the first equation.x - 3y = -6 2x - 7y = 10 Answers: 1/3 x +2 -1/3 x + 6 -x + 2 -1/3 x + 2

OpenStudy (anonymous):

d

OpenStudy (anonymous):

x - 3y = -6 subtract x from both sides then divide by -3 what do you get ?

OpenStudy (anonymous):

(x,y) = (-72,-22)

OpenStudy (anonymous):

(i got D on my paper...)

OpenStudy (anonymous):

Hey you got good equations now...that's what its D...

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

Actually the answer is A

OpenStudy (anonymous):

x - 3y = -6 subtract x from both sides then divide by -3 what do you get : -3y = -6 -x y = 2 + (1/3)x

OpenStudy (anonymous):

answer is A indeed ..

OpenStudy (anonymous):

Yep its A..

OpenStudy (anonymous):

(-72,-22) satisfy A, not D...Sorry:(

OpenStudy (anonymous):

The answer I got was A when I checked it, not D...?

OpenStudy (anonymous):

A